OpenStudy (anonymous):

The distance travelled in 2 nd second is 12m , and 20 m in 4 th second what will the distance covered in 4 seconds after the 5 th second?

OpenStudy (anonymous):

v=d2-d1/t2-t1=20-12/4-2=4ms^-1 so now 4 seconds after 5th second wil be 4*4 seconds=16m this distance is travelled after 5th second then distance travelled at 5th second is 20+4m=24m so answer wil be distance at 5th second +distance after 4 seconds 24+!6=40

OpenStudy (anonymous):

@Yahoo!

OpenStudy (anonymous):

Dn = U + a/2 (2n-1)

OpenStudy (anonymous):

@theyatin

OpenStudy (anonymous):

well @yahoo you were teaching or asking?? sorry i f i am wrong this is all i have. . .

OpenStudy (anonymous):

U r correct lol

OpenStudy (anonymous):

40m is the answer

OpenStudy (anonymous):

@Yahoo! well i dont always follow topology ;)

OpenStudy (anonymous):

thats ok.............thxxxx a lot